OpenStudy (anonymous):

if (x+2)(x-6)=0 then we say either x+4=0 or x-6=0 why not we say x+4=0 and x-6=0?

OpenStudy (anonymous):

Well x is either -4 or it's 6, it can't be both...

OpenStudy (anonymous):

That's... yeah... pretty much the sum of it. x can only have a distinct value.

OpenStudy (anonymous):

if (x+2)(x-6)=0 then we can not say either x+4=0 or x-6=0

OpenStudy (anonymous):

pretend that x is like time. It is either 4 minutes ago or 6 minutes in the future, it cannot be both at the same time not even in Back to the Future

OpenStudy (anonymous):

Ha! FFM! We can say that x+2 = 0 or x-6 = 0 :-P

OpenStudy (anonymous):

Yep!

OpenStudy (anonymous):

ok when we have xy=0 then why we say either x=0 or y=0 why not we say x=0 and y=0

OpenStudy (anonymous):

Hahaha you were the only one to notice.

OpenStudy (anonymous):

It's an inclusive "or".

OpenStudy (anonymous):

If it's xy = 0 then x and y COULD both be 0, or one could be 0.

OpenStudy (anonymous):

its either

OpenStudy (anonymous):

You can't work anything out from xy = 0. If you're given that as an equation, the absolute best you can calculate is that either one or both x and y are equal to 0. Beyond that only speculation is possible, as if one variable is 0 the product is 0 and the other could therefore be anything.

OpenStudy (anonymous):

(I mean literally xy = 0. Obviously you can work out {equation containing x and y} = 0).

OpenStudy (anonymous):

if (x+4)(x-6)=0 then we say either x+4=0 or x-6=0 The constraint "or" is to avoid division by zero.

OpenStudy (anonymous):

are all of u sure that if xy=0 then both x and y can be zero. 100% sure?

OpenStudy (anonymous):

It is like saying I will go get groceries on Tuesday or Wednesday, you need to do it one of the days, but you may go both days as well.

OpenStudy (anonymous):

Yes :)

OpenStudy (anonymous):

They can both be zero, absolutely, but they do not have to both be zero and in most math problems the are not both zero.

OpenStudy (anonymous):

\( xy=0 \) If both are zero, then \( x = \frac{0}{y} \) but y is zero too, Disaster !!!

OpenStudy (anonymous):

FFM I dunno if this is the thread to bring that up :)

OpenStudy (anonymous):

have any of u read vector space. in which i have read if v is a vector and k is a constant/scalar then if kv=0 and v=0 then v is linearly dependent bcz k would be non zero.

OpenStudy (anonymous):

That is not correct, regarding vector spaces.

OpenStudy (anonymous):

But if we have xy = 0, we cannot assume x = 0/y is a viable option. If we have xy = 0, then they both may be zero.

OpenStudy (anonymous):

>"But if we have xy = 0, we cannot assume x = 0/y is a viable option" why ? That's an equation, division by \( y \) on both sides is perfectly valid.

OpenStudy (anonymous):

No it's not. Division by zero is never valid.

OpenStudy (anonymous):

Yes that's why we can't have both zero.

OpenStudy (zarkon):

who is to say that y is a function of x?

OpenStudy (anonymous):

what is the logic of ur answer. anything is correct for calculus is false for vector spaces.

OpenStudy (anonymous):

That's why the "or"

OpenStudy (anonymous):

That's not true at all.

OpenStudy (zarkon):

i define x=0 and y=0 then xy=0.....both zero

OpenStudy (anonymous):

You may only divide if you make the assumption that y is not zero.

OpenStudy (anonymous):

lol, Zarkon that's alright then, I assumed y is the dependent variable ;)

OpenStudy (anonymous):

I don't understand what that has to do with anything... two quantities multiplied together equals zero. Either the first is zero, the second is zero, or both are.

OpenStudy (anonymous):

It depends on the definition on X = (x+4) and Y=(x-4), and (x+4)(x-4)=0 then we can't have both zero, as they have same independent variable

hero (hero):

in order for ab = 0, only one value needs to be zero. They can't both be zero because a doesn't equal b. They have to be two different numbers.

OpenStudy (anonymous):

Just because we have multiplication, it does not mean that we can divide. This is the reason we get extraneous solutions in algebra. If we have a = 5 and x(a - 5) = 0 we cannot divide both sides by a-5.

OpenStudy (anonymous):

but if x and y are not defined then?

OpenStudy (anonymous):

what do you mean be not defined ? They have to defined.

OpenStudy (anonymous):

Obviously if y = (x+2) both can't be zero but we resolved that question like four posts in. In general, xy = 0 implies x = 0, y = 0, or both x and y = 0.

OpenStudy (anonymous):

In general, xy = 0 implies x = 0, y = 0, or both x and y = 0. -- No, the both part depends on the definition of x and y.

OpenStudy (anonymous):

ab = 144 implies that a and b are different? Abe is 12 and Barney is 12, they are twin brothers, who happen to be 12, they cannot be the same age?

OpenStudy (zarkon):

as for your vector space question...are you asking if the the set consisting of only the zero vector linearly dependent?

OpenStudy (anonymous):

In general. In general. In general.

OpenStudy (anonymous):

If they are not defined, then they are unknown and they may BOTH be zero

OpenStudy (anonymous):

In general I always assume y is a function of x ;)

hero (hero):

a = 3 b = 0 3*0 = 0 a = 0, b = 4 0*4 = 0 When dealing with factorable quadratics, a and b will never both be zero.

OpenStudy (anonymous):

haha FFM have you never worked with two variables, then? And Hero, yeah, but that's not what I am talking about.

OpenStudy (anonymous):

yes zarkon i want to ask the same

OpenStudy (anonymous):

(0)(0)=0

OpenStudy (anonymous):

For a^2 - b^2 = 0 One solution is a = 4 and b = 4

hero (hero):

JeMurray, I was only talking to Rosy, not you.

OpenStudy (zarkon):

the set \[S=\{\vec{0}\}\] is a linearly dependent set

OpenStudy (anonymous):

Duly noted Hero, sorry :)

OpenStudy (anonymous):

they can be both equal to 0

OpenStudy (anonymous):

how it is linearly dependent zarkon?

OpenStudy (zarkon):

it is dependent because you can write a linear combination equal to zero in a nontrivial way

OpenStudy (anonymous):

@zarkon that's what i want to ask if we write it in non-trivial way then it means scalar is non zero i.e if kv=0 and v=0 then k must be non zero. means in kv=0 k and v both can't be both zero but in xy=0 all of you have said that both can be zero what is this?????????????????????? please answer me soon.........

OpenStudy (anonymous):

Let me clarify what I think you mean when you talk about vector spaces. Let V be a vector space containing three vectors, \[ V = \{\vec{v}_1, \vec{v}_2, \vec{v}_3 \}\] The vectors \[\vec{v}_1,\vec{v}_2,\text{ and } \vec{v}_3\] are linearly dependent if you can write \[a\cdot \vec{v}_1 + b\cdot \vec{v_2} + c\cdot \vec{v_3} = \vec{0}\] where a, b, and c are not all zero. Zarkon said that the collection \[V = \{\vec{0}\} \] is linearly dependent. This is of course true, because \[ a \cdot \vec{0} = \vec{0} \] for any constant number a.

OpenStudy (anonymous):

yes but it can be linearly dependent when a is not equal to zero and question still remain the same why it can not be zero why?????????????

OpenStudy (anonymous):

There is no mathematical reason that a could not be zero. We simply don't allow a to be zero when we're talking about linear dependence.

OpenStudy (zarkon):

a=0 works (\(0\vec{0}=\vec{0}\)), but so does a=5 or a=-7 all the definition says (for linear dependent) is that there is a way to write the linear combination so that not all the coefficients are zero, but the L.C. is the zero vector

OpenStudy (anonymous):

i think you do not understand my que. i am just asking that in linear dependence we say if v=0 then kv=0 is linear dependent bcz k is non zero on the other hand in xy=0 we say both x and y can be zero. isn't it strange???

OpenStudy (zarkon):

S={v} , (v=\(\vec{0}\)) is linearly dependent because it is possible to write kv=\(\vec{0}\) where k is not zero. you could also let k=0 then it is still true that kv=\(\vec{0}\)